OpenStudy (toxicsugar22):

For the following reaction, 5.28 grams of perchloric acid (HClO4) are mixed with excess tetraphosphorus decaoxide. The reaction yields 1.33 grams of phosphoric acid. 12 HClO4 (aq) + P4O10 (s) = 4 H3PO4 (aq) + 6 Cl2O7 (g) (1) What is the theoretical yield of phosphoric acid ? grams (2) What is the percent yield for this reaction ? %

OpenStudy (ghazi):

molar mass of HClO4 = 100 molar mass of P4O10 = 284 molar mass of H3PO4 = 98 molar mass of CL2O7= 182 now you can do it, its simple stoichiometry

OpenStudy (ghazi):

5.28gms of HClO4 means moles = \[n= \frac{ 5.28 }{ 100 }= 0.0528 \]

OpenStudy (ghazi):

hold on now, i am doing it

OpenStudy (ghazi):

now according to equation 12 moles of HClO4 yields 4 moles of H3PO4 so 1 mole of HCLO4 WILL GIVE 4/12 MOLES OF H3PO4 OR 1/3 MOLES OF H3PO4 therefore 0.0528 moles of HClO4 will yeild 0.0528*1/3 moles of H3PO4 = 0.0176 moles , this is your first part

OpenStudy (ghazi):

well thats a nasty calculation i am tired of writing :(

OpenStudy (ghazi):

25.18%

OpenStudy (ghazi):

what is it ?

OpenStudy (toxicsugar22):

For the following reaction, 3.88 grams of glucose (C6H12O6) are mixed with excess oxygen gas. The reaction yields 4.56 grams of carbon dioxide. C6H12O6 (s) + 6 O2 (g) = 6 CO2 (g) + 6 H2O (l) (1) What is the theoretical yield of carbon dioxide ? grams (2) What is the percent yield for this reaction ? %

OpenStudy (ghazi):

1 mole of glucose gives 6 moles of CO2 3.88 gms of glucose = 0.021555 moles of glucose so 0.02155 moles of glucose will give 0.02155*6 = 0.1293 moles of co2 , first part

OpenStudy (toxicsugar22):

ok

OpenStudy (ghazi):

cool :)

OpenStudy (toxicsugar22):

then for the second part it would be .1293 times 100

OpenStudy (ghazi):

umm i dont think so you have to do a minor calculation and my calculation says a weird output because i am sleepy lol and its seven in the morning so i gotta go :( and i think answer to your second part is more than 1, that is around 1.207

OpenStudy (anonymous):

WHAT DO YOU WANT ME TO DO

OpenStudy (anonymous):

OK GONE